# Rethinking Bottlenecks in Safety Fine-Tuning of Vision Language Models

### v1

@@ 3.2 Construction Pipeline
Models ASR($\downarrow$) HR($\downarrow$) RSR($\uparrow$) RR($\uparrow$) ASR($\downarrow$) HR($\downarrow$) RSR($\uparrow$) RR($\uparrow$) ASR($\downarrow$) HR($\downarrow$) RSR($\uparrow$) RR($\uparrow$)
@@
3.3 MIRage
Similar to previous SFT methods, we added 500 general QA samples from training set of M4-Instruct (Li et al., 2024b) to ensure the model’s ability to follow instructions remains unaffected. Notably, in the final training set of 4.5k samples, only 11% are general samples, much lower than the 33% used in Textual SFT (Hu et al., 2024) and VLGuard (Zong et al., 2024). We provide detailed SFT information in Section A.1.
4 Experiments
4.1 Setup
Training Details.
Our main experiments focus on applying MIRage to InternVL2.5-8B (Chen et al., 2024c), with results on additional models presented in the Appendix B. We fully fine-tune models on 8*A100-80G GPU with a global batch size of 128, requiring only 30 minutes of training time. More training details also are provided in Section A.1.
Baselines.
We evaluate 13 VLMs with robust multi-image understanding capabilities, including open-sourced models: LLaVA-NeXT-Interleave (Li et al., 2024b), LLaVA-OV-72b-Chat (Li et al., 2024a), Mantis-SIGLIP (Jiang et al., 2024), Idefics3-8B (Laurençon et al., 2024), Phi-3.5-Vision (Abdin et al., 2024), MiniCPM-V 2.6 (Yao et al., 2024), DeepSeek-VL2 (Wu et al., 2024), Qwen2-VL-7B&72B-Instruct (Wang et al., 2024b), and InternVL2.5-8B&78B (Chen et al., 2024c); close-source API model: GPT-4o (Achiam et al., 2023), and Gemini-1.5-Pro (Team et al., 2024). Additionally, we also evaluate some safety fine-tuning baselines, including Textual SFT (Hu et al., 2024) and VLGuard-R on InternVL2.5-8B. We deploy and inference all models on vLLM (Kwon et al., 2023).
Evaluation on MIS Test Set.
We use GPT-4o (Achiam et al., 2023) to classify all responses into four categories: Unsafe, for responses that present safety risks based on harmful instructions; Safe with Reasoning, for answers that identify content in the images and logically deduce potential harmful intent with a warning; Safe with Refusal, for brief refusal responses; and Hallucination, for responses that are irrelevant or incomplete due to the model’s failure to understand the samples. Based on these four categories, we use Attack Success Rate (ASR), Reasoning Success Rate (RSR), Refusal Rate (RR), and Hallucination Rate (HR) as metrics to evaluate the model’s safety capabilities. More detailed evaluation can be found in the Section A.2.1.
Evaluation Benchmarks on MIRage.
We evaluate our MIRage from two perspectives: Harmlessness and Helpfulness. Building on the settings from Sec. 2.1, we conduct additional experiments on general benchmarks such as MuirBench (Wang et al., 2024a), and Q-Bench (Wu et al., 2023). Furthermore, we introduce more safety tasks, including MM-SafetyBench (Liu et al., 2025), Jailbreak-V mini-set (Luo et al., 2024), and the challenging SIUO task (Wang et al., 2024c). Detailed settings and benchmark description can be found in Section A.2.
4.2 Main Result
In this subsection, we evaluate and analyze the performance of open-sourced VLMs, closed-source API models, and MIRage on MIS test set. We report the main result in Table 4.
Most VLMs Lack Multi-Image Safety Awareness Entirely.
@@ 4.2 Main Result
Recalling results in Table 1, recent VLMs and fine-tuning methods have demonstrated strong safety capabilities in single-image settings. However, as shown in Table 4, we find that the safety awareness aligned in these settings does not generalize well to the multi-image setting. Especially for open-source models, except for Phi3.5-Vision, other models exhibit ASR of around 90% across different test sets. Interestingly, most VLMs exhibit a lower ASR on the MIS-easy set than the MIS-hard set. We attribute this to the models’ safety and reasoning abilities in multi-image settings are too weak to understand the harmful intent in the challenging samples, leading them to produce irrelevant answers based on only one of the images. For the low ASR of Phi3.5-Vision, we found that the technical report emphasizes multimodal safety alignment. However, our evaluation revealed many incomplete and hallucinated responses, leading to higher HR.
Close-source API Models also Easy to Jailbreak.
Even API models like GPT-4o show an ASR of 46.21% and 65.29% for the MIS-easy and MIS-hard sets, respectively. Similar to the findings of Hu et al. (2024), Gemini-1.5-pro exhibits stronger safety capabilities compared to other models, achieving an ASR of approximately 40% on both MIS-easy and MIS-hard sets. Additionally, we observe that Gemini often provides safe responses through reasoning rather than direct refusals, contributing to its lower ASR. However, even strong API models remain susceptible to defense jailbreak in multi-image settings.
@@
Table 5: Comparison of safety fine-tuning methods on general ability benchmarks. The Average term represents the average accuracy across these 5 tasks. Baselines with contain 500 general data in MIRage, while indicates MIRage without the 500 general data.
General Data
Q-Bench SI
MMStar SI
MMMU SI&MI
MuirBench MI
MMT MI
Average
Models (%) Exactly Match($\uparrow$)
@@
Synthetic Images Are More Easily to Jailbreak than Real Images.
Interestingly, we observe that the ASR of the MIS-real set is slightly lower than that of the MIS-easy and MIS-hard set. We hypothesize that this is because the retrieved images are generally simpler, containing only the generated objects. Moreover, the distribution of real images is likely closer to the model’s training data, which may lead to the model making more accurate safety inferences on these images. However, as the quality of T2I-generated images improves, some safety benchmarks are also leveraging synthetic images (Liu et al., 2025; Hu et al., 2024). This finding highlights the importance of improving model defenses against synthetic images in the future.
MIRage Significantly Increase Multi-Image Safety of Base Model.
We present the results of our MIRage on InternVL2.5-8B in Table 4. After fine-tuning with MIRage, the nearly 0 values of ASR and HR indicate a significant improvement in safety capability. By observing the results of RSR and RR, we find that, as expected, the introduction of reasoning logic into the SFT labels causes the fine-tuned model to rely more on visual reasoning to produce safer responses. We apply MIRage on more VLMs and validate that vanilla CoT cannot solve MIS tasks in Appendix B.
Table 6: Comparison of safety-related tasks across different safety fine-tuning methods. MSS and $M$-Safe represent MSSBench and $M$-SafetyBench, respectively. Among these benchmarks, FigStep and $M$-Safety involve OCR-based attacks; JailbreakV is a multimodal jailbreak dataset, with its mini-set used here; while SIUO and MSS are challenging safety tasks, designed to trigger unsafe responses from models using benign image-text pairs.
FigStep
$M$-Safety
JailbreakV
SIUO
MSS Safe
MSS Unsafe
[…]
InternVL2.5-8B 38.80 15.58 18.57 24.85 99.67 3.00
+ Textual SFT 30.60 2.54 6.37 20.61 99.33 1.00
+ VLGuard-R 0.60 0.66 3.67 64.23 88.33 35.44
+ MIRage 0.60 0.54 3.21 71.26 87.67 40.00
4.3 Analysis about MIRage
In this subsection, we conduct more experiments on helpfulness and harmlessness to validate the superiority of MIRage.
@@ 4.3 Analysis about MIRage
[…]
Figure 4: Comparison on our MIS test set. Top: ASR $(\downarrow)$( ↓ ) of different methods. Bottom: RSR $(\uparrow)$( ↑ ) of different methods.
Generalizable Safety through Enhanced Reasoning Ability.
Previously, we validated in Sec. 4.2 that MIRage enhances the model’s multi-image safety capability. However, it is also important to investigate its generalization and the underlying reasons for the performance improvement. In Table 6 and Figure 4, we compare MIRage with other baselines. Interestingly, we find that Textual SFT is not able to handle challenging safety tasks with benign text instructions. Additionally, VLGuard-R shows similar safety performance on relatively easy tasks with MIRage. However, in tasks like SIUO, MSSBench Unsafe, and our MIS test set, which require more visual reasoning to identify potential unsafe intent and provide safe responses, MIRage outperforms the baselines. This suggests that MIRage achieves safety through improved visual reasoning ability.
@@ 5 Related Works
With the rapid advancement of VLM capabilities (Achiam et al., 2023; Chen et al., 2024c), it becomes increasingly crucial to mitigate the risks associated with unsafe instructions. Similarly to aligning language models, safeguarding VLMs always involves collecting external data and feedback related to safety. Zhang et al. (2024b) utilizes preference optimization methods such as RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2024), steering model to more safe distribution by constructing 90k safety preference data SPA-VL. Alternatively, Li et al. (2024d); Zong et al. (2024) adopt more resource-efficient SFT for safety feedback, introducing RTVLM and VLGuard, each containing 5.2k and 2k safety-related question-answer pairs, respectively. Although these multimodal fine-tuning methods effectively provide guardrails for VLMs, Guo et al. (2024b) points out that they tend to exhibit over-prudence when facing neutral data. Wang et al. (2024c) found that safety ability achieved by unlearning in the textual space can generalize to multimodal situations with minimal impact on general performance. However, Hu et al. (2024) highlighted that the safety acquired in textual space only applies to cases where harmful visual information is leaked through text instructions, limiting its generalizability. Therefore, fine-tuning methods still face bottlenecks in achieving comprehensive safety alignment.
@@
Researchers have made significant strides in evaluating the multimodal safety capabilities of models (Gong et al., 2023; Li et al., 2024d; Zhang et al., 2024b; Liu et al., 2025; Hu et al., 2024). For example, VLSafe (Chen et al., 2024b) and SPA-VL (Zhang et al., 2024b) pair harmful instructions with related images to create multimodal safety settings. Additionally, Hu et al. (2024) introduced visual safety information leakage in VLSBench, using images to express unsafe intent. Similarly, FigStep (Gong et al., 2023) and MM-SafetyBench (Liu et al., 2025) use OCR-based attacks via the vision modality. Wang et al. (2024c) further demonstrates that unsafe content can be triggered using benign text-image pairs and introduces SIUO, while Zhou et al. (2024) presents Multimodal Situational Safety, where different safe image scenarios are paired with the same benign query to create both safe and unsafe situations. However, all these benchmarks are based on single-image settings. As VLMs rapidly advance in complex visual capabilities, particularly multi-image understanding, evaluating their safety in multi-image contexts has become crucial. To the best of our knowledge, we introduce MIS, the first multi-image safety dataset, bridging the gap in evaluating VLMs’ safety abilities in multi-image settings.
6 Conclusion
In this paper, we identify bottlenecks of existing safety fine-tuning methods, which fails to be adequetely effective or lead to over-prudence behavior. To address this challenge, we introduce MIS, the first dataset for improving and evaluating VLMs’ safety-related performance on both visual perception and visual reasoning. Alongside the training set, we propose MIRage, a novel safety fine-tuning paradigm that improves both helpfulness and safety by enhancing reasoning ability. Our results reveal significant vulnerabilities in current VLMs’ multi-image safety and demonstrate that safety fine-tuning with multi-image data can not only reduce ASR on MIS, but also generalize across various safety tasks, while without trade-offs on general performance, highlighting the potential of multi-image data for safety fine-tuning.
7 Impact Statement
@@
Appendix A Detailed Experiment Settings
A.1 Training Details
Table 7: Detail training parameters of Textual SFT on SafeRLHF, Single-Image Multimodal SFT on VLGuard-P and VLguard-R, and our Multi-Image Multimodal SFT MIRage. Methods with are added additional 500 general QA samples in the MIS training set. Mirage with represents that without 500 general QA samples in fine-tuning data.
@@ A.1 Training Details
We apply different fine-tuning strategies on two powerful open-sourced VLMs: Qwen2-VL-7B-Instruct (Wang et al., 2024b) and InternVL2.5-8B (Chen et al., 2024c). For Qwen2-VL, we fine-tune using LlamaFactory (Zheng et al., 2024), as recommended in their official GitHub repository. Notably, for Textual SFT, we only fine-tune the LLM backbone weights, whereas for Multimodal SFT, we fine-tune both the LLM backbone and projector parameters, with all experiments performed using full parameter fine-tuning. For InternVL2.5, we use the fine-tuning scripts provided in their official documentation [^1] https://internvl.readthedocs.io/en/latest/internvl2.5/finetune.html. All experiments are conducted on 8*A100-80G GPUs.
@@
For Textual SFT, we follow the settings of (Hu et al., 2024), sampling 2k harmful and 1k benign samples from SafeRLHF (Ji et al., 2024). For the harmful input data, we use Llama3-8B-Instruct (Inan et al., 2023) to generate safe responses, while the safe samples use the original responses from the dataset. Since the data is in the text domain, following (Hu et al., 2024), we match each sample with an all-black image during fine-tuning. It is worth noting that for Qwen2-VL, we directly use the open-sourced fine-tuned model [^2] https://huggingface.co/Foreshhh/Qwen2-VL-7B-SafeRLHF from (Hu et al., 2024), while for the recent InternVL2.5, the detailed training parameters are reported in Table 7.
@@
Given that we cannot access the training data for Qwen2-VL and InternVL2.5, we can only conduct experiments on VLGuard Posthoc Fine-Tuning (Zong et al., 2024), which we refer to as VLGuard-P. This includes 2k multimodal single-image safety-related data and 1k benign samples. As for VLGuard-R, mentioned in Sec. 2, we designed a prompt that guides InternVL2.5-78B to first analyze the input content and provide reasons for its potential unsafe nature, ultimately generating harmless responses. This process was used to construct SFT labels for 2k unsafe inputs. For the 1k safe samples, we directly use the original responses from VLGuard. The specific prompt is reported in Figure. As with Textual SFT, for Qwen2-VL, we use the open-sourced VLGuard-P [^3] https://huggingface.co/Foreshhh/Qwen2-VL-7B-VLGuard from (Hu et al., 2024). The training details for VLGuard-R and InternVL2.5 are provided in Table 7. Additionally, the detailed SFT label construction prompt is shown in Sec. E.2.
@@
With our proposed MIS training set, which contains 4k safety-related multi-image samples, we introduce MIRage. Given that safety risks in MIS arise from text instructions linking objects across two images, we design a prompt to guide InternVL2.5-78B to respond based on safety reasoning logic. Specifically, as shown in the figure, the model is first instructed to identify the elements in both images, then reason about the potential risks created by these elements under the text instruction, and finally provide a safe response. Building on this, we additionally include 500 general QA samples to ensure that the model’s instruction-following capability remains unaffected. The detailed training parameters are provided in Table 7. Additionally, the detailed SFT label construction prompt is shown in Sec. E.2.
A.2 Evaluation Details
A.2.1 Evaluation on MIS

### v3

@@ 4.1 Main Result
[…]
@@
Qwen2-VL-7B+MIRage 1.67 0.00 97.61 0.72 1.76 0.00 98.24 0.00 1.00 0.00 98.00 1.00
MiniCPM-V2.6+MIRage 1.91 0.00 96.90 1.19 1.57 0.00 98.43 0.00 1.00 0.00 98.00 1.00
LLaVA-OV-7B+MIRage 1.55 0.00 97.26 1.19 0.78 0.00 99.22 0.00 2.00 0.00 97.00 1.00
@@
Most VLMs Lack Multi-Image Safety Awareness.
@@
As shown in Table 1, recent VLMs and fine-tuning methods demonstrate strong safety capabilities in single-image settings. However, Table 4 reveals that such alignment does not generalize well to multi-image scenarios. Open-source models, except for Phi3.5-Vision, show ASR rates around 90% across test sets, and interestingly, most exhibit lower ASR on MIS-easy than MIS-hard. We attribute this to their weak safety and reasoning abilities in multi-image settings, which prevent them from detecting harmful intent in challenging samples and instead lead to irrelevant answers based on only one image. Phi3.5-Vision achieves lower ASR by emphasizing multimodal safety alignment, but our evaluation reveals many incomplete or hallucinated responses, resulting in higher HR. Even closed-source API models are not immune: GPT-4o shows ASR of 46.21% and 65.29% on easy and hard, respectively, while Gemini-1.5-pro performs better with ASR around 40% on both sets. Additionally, we observe that Gemini often provides safe responses through reasoning rather than direct refusals, contributing to its lower ASR. However, even strong API models remain susceptible to defense jailbreak in multi-image settings.
Synthetic Images Are Easier to Jailbreak than Real Ones.
We observe that the ASR of the MIS-real is slightly lower than that of the MIS-easy and MIS-hard. We hypothesize that this is because the retrieved images are generally simpler, containing only the generated objects. Moreover, real images are likely closer to the model’s training distribution, enabling more accurate safety inferences. Yet as T2I-generated images improve, some safety benchmarks increasingly adopt synthetic data (Miao et al., 2025), underscoring the need to strengthen model defenses against such images.
Figure 4: Comparison of ASR $(\downarrow)$ and RSR $(\uparrow)$ of different methods on MIS test set.
MIRage Significantly Increases Safety Ability in Both Single and Multi-Image Settings.
We present the results of MIRage on InternVL2.5-8B in Table 4. After fine-tuning with MIRage, the near-zero ASR and HR values indicate substantial safety improvements. Examining RSR and RR further shows that introducing reasoning logic into the SFT labels encourages the fine-tuned model to rely more on visual reasoning for safer responses. We extend MIRage to additional VLMs and confirm in Appendix C that a vanilla CoT prompt alone cannot solve MIS tasks. Moreover, Table 5 and Fig. 4 compare MIRage with other baselines. Interestingly, Textual SFT struggles with challenging safety tasks involving benign instructions, while VLGuard-R achieves comparable performance to MIRage on relatively easy cases. However, in tasks such as SIUO, MSSBench Unsafe, and our MIS benchmark, where visual reasoning is critical for detecting unsafe intent, MIRage consistently outperforms baselines. These results suggest that MIRage enhances safety primarily through improved visual reasoning.
Table 5: Comparison of safety-related tasks across different safety fine-tuning methods. MSS and $M$-Safe represent MSSBench and $M$-SafetyBench, respectively.
FigStep $M$-Safety JailbreakV SIUO MSS Safe MSS Unsafe
Models ASR $\downarrow$ Safe $\uparrow$ Acc $\uparrow$
InternVL2.5-8B 38.80 15.58 18.57 24.85 99.67 3.00
+ Textual SFT 30.60 2.54 6.37 20.61 99.33 1.00
+ VLGuard-R 0.60 0.66 3.67 64.23 88.33 35.44
+ MIRage 0.60 0.54 3.21 71.26 87.67 40.00
4.2 More Discussion about MIRage
Generalizable Safety through Enhanced Reasoning Ability.
Figure 5: Safety capabilities acquired by MIRage generalize to unseen categories.
We validated in Sec. 4.1 that MIRage enhances the model’s safety capability. However, it is also important to investigate its generalization and the underlying reasons for the performance improvement. To investigate whether the safety capabilities acquired through MIRage can generalize to unseen safety categories, we follow the VLGuard setup for evaluation. Specifically, we remove the Privacy and Self-Harm categories from the MIS training set. During testing, we evaluate the model on the corresponding categories in the MIS test set, VLSBench, and the Privacy category in MM-SafetyBench. As shown in Fig. 5, the results demonstrate that the safety capabilities learned by MIRage generalize well to previously unseen safety categories. We conduct more ablation study and analysis in Appendix C.
@@ 4.2 More Discussion about MIRage
Table 6: Comparison of safety fine-tuning methods on general ability benchmarks. Average denotes the mean accuracy across the five tasks. Baselines marked with include 500 general samples from MIRage, incorporates 6000 additional general samples randomly sampled from other sources, and denotes MIRage without the 500 general samples.
General Data Q-Bench (SI) MMStar (SI) MMMU (SI&MI) MuirBench (MI) MMT (MI) Average
Models (%) Exact Match ($\uparrow$)
@@
Previous studies suggest that incorporating helpfulness data is crucial to avoid overly conservative safety behaviors (Zong et al., 2024). For example, VLGuard augments its training set with 5k general samples from LLaVA-v1.5 (Liu et al., 2024a) and 1k from VLGuard itself, which is three times the amount of its unsafe data, to preserve generalization. Textual SFT includes 1k general-safe samples, amounting to 0.5 times its unsafe data. Despite this, as shown in Table 1, both methods show limited general performance. In contrast, our MIRage uses only 500 general QA samples, which represent just 11% of its unsafe data. For a fair comparison, we also add these 500 samples to VLGuard and Textual SFT, resulting in general-to-unsafe data ratios of 33% and 42%, respectively. As shown in Table 6, MIRage achieves the best helpfulness performance with the smallest amount of general data, even slightly outperforming the base model. This highlights the strong visual understanding and reasoning capabilities gained through multi-image training. Although VLGuard-R introduces reasoning labels, its simple single-image inputs limit further improvement. Furthermore, adding 5k extra general samples to VLGuard does not enhance its generalization ability; it only mitigates the model’s over-prudence behavior. More results are provided in Appendix C.5.
@@ 5 Related Works
With the rapid advancement of VLM capabilities (Achiam et al., 2023; Chen et al., 2024c), it becomes increasingly crucial to mitigate the risks associated with unsafe instructions. Similarly to aligning language models, safeguarding VLMs always involves collecting external data and feedback related to safety. Zhang et al. (2024b) utilizes preference optimization methods such as RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2024), steering model to safer distribution by constructing 90k safety preference data SPA-VL. Alternatively, Li et al. (2024d); Zong et al. (2024) adopt more resource-efficient SFT for safety feedback, introducing RTVLM and VLGuard, each containing 5.2k and 2k safety-related question-answer pairs, respectively. Although these multimodal fine-tuning methods effectively provide guardrails for VLMs, Guo et al. (2024b) points out that they tend to exhibit over-prudence when facing neutral data. Wang et al. (2024c) found that safety ability achieved by unlearning in the textual space can generalize to multimodal situations with minimal impact on general performance. However, Hu et al. (2024) highlighted that the safety acquired in textual space only applies to cases where harmful visual information is leaked through text instructions, limiting its generalizability. Therefore, fine-tuning methods still face bottlenecks in achieving comprehensive safety alignment.
@@
Researchers have made significant strides in evaluating the multimodal safety capabilities of models (Gong et al., 2023; Li et al., 2024d; Zhang et al., 2024b; Liu et al., 2025; Hu et al., 2024; Ziqi et al., 2025). For example, VLSafe (Chen et al., 2024b) and SPA-VL (Zhang et al., 2024b) pair harmful instructions with related images to create multimodal safety settings. Additionally, Hu et al. (2024) introduced visual safety information leakage in VLSBench, using images to express unsafe intent. Similarly, FigStep (Gong et al., 2023) and MM-SafetyBench (Liu et al., 2025) use OCR-based attacks via the vision modality. Wang et al. (2024c) further demonstrates that unsafe content can be triggered using benign text-image pairs and introduces SIUO, while (Zhou et al., 2024) presents Multimodal Situational Safety, where different safe image scenarios are paired with the same benign query to create both safe and unsafe situations. However, all these benchmarks are based on single-image settings. As VLMs rapidly advance in complex visual capabilities, particularly multi-image understanding, evaluating their safety in multi-image contexts has become crucial. To the best of our knowledge, we introduce MIS, the first multi-image safety dataset, bridging the gap in evaluating VLMs’ safety abilities in multi-image settings.
6 Conclusion and Limitation
In this paper, we identify bottlenecks of existing safety fine-tuning methods, which fails to be adequetely effective or lead to over-prudence behavior. To address this challenge, we introduce MIS, the first dataset for improving and evaluating VLMs’ safety-related performance on both visual perception and visual reasoning. Alongside the training set, we propose MIRage, a novel safety fine-tuning paradigm that improves both helpfulness and safety by enhancing reasoning ability. Our results reveal significant vulnerabilities in current VLMs’ multi-image safety and demonstrate that safety fine-tuning with multi-image data can not only reduce ASR on MIS, but also generalize across various safety tasks, while without trade-offs on general performance, highlighting the potential of multi-image data for safety fine-tuning. Our MIRage framework explores a simple approach to constructing Safety CoT labels that incorporate visual perception, visual reasoning, and safe response generation. In future work, potential directions include designing more complex reasoning labels or leveraging reinforcement learning to help models acquire stronger safety reasoning capabilities.
Ethics statement
@@ 6 Conclusion and Limitation
Reproducibility statement
The methods introduced in this paper are described in detail in Sec. 3, with implementation details provided in Appendix B. Their code implementations are included as anonymous, downloadable source files in the supplementary materials.
Acknowledgement
This work was supported by Shanghai Artificial Intelligence Laboratory.
Appendix A LLM Usage Statement
The construction of the codebase partially relied on AI assistant for debugging, and AI assistant polished the writing of this paper.
Appendix B Detailed Experiment Settings
B.1 Detailed Baselines
We evaluate 14 VLMs with robust multi-image understanding capabilities, including 11 open-sourced models, including: LLaVA-NeXT-Interleave (Li et al., 2024b), LLaVA-OV-7b, LLaVA-OV-72b-Chat (Li et al., 2024a), Mantis-SIGLIP (Jiang et al., 2024), Idefics3-8B (Laurençon et al., 2024), Phi-3.5-Vision (Abdin et al., 2024), MiniCPM-V 2.6 (Yao et al., 2024), DeepSeek-VL2 (Wu et al., 2024), Qwen2-VL-7B&72B-Instruct (Wang et al., 2024b), and InternVL2.5-8B&78B (Chen et al., 2024c); and 2 closed-source API models: GPT-4o (Achiam et al., 2023), and Gemini-1.5-Pro (Team et al., 2024). Additionally, we also evaluate some safety fine-tuning baselines, including Textual SFT and VLGuard-R on 4 open-sourced VLMs. We deploy and inference all models on vLLM (Kwon et al., 2023).
B.2 Training Details
Table 7: Detail training parameters of Textual SFT, Single-Image Multimodal SFT on VLGuard-P and VLguard-R, and our Multi-Image Multimodal SFT MIRage. Methods Baselines marked with include 500 general samples from MIRage, incorporates 6000 additional general samples randomly sampled from other sources, and denotes MIRage without the 500 general samples.
@@ B.2 Training Details
MiniCPM-V 2.6 Textual SFT 33.33 2e-5 0.03 2
VLGuard-R 33.33 5e-6 0.03 3
MIRage 11.11 5e-6 0.03 3
LLaVA-OV-7B Textual SFT 33.33 2e-5 0.03 3
VLGuard-R 33.33 5e-6 0.03 3
MIRage 11.11 5e-6 0.03 3
@@
We apply different fine-tuning strategies on four powerful open-sourced VLMs: Qwen2-VL-7B-Instruct Wang et al. (2024b), InternVL2.5-8B Chen et al. (2024c), MiniCPM-V 2.6 (Yao et al., 2024), and LLaVA-OV-7B (Li et al., 2024a). For Qwen2-VL and MiniCPM-V 2.6, we fine-tune using LlamaFactory Zheng et al. (2024), as recommended in their official GitHub repository. Notably, for Textual SFT, we only fine-tune the LLM backbone weights, whereas for Multimodal SFT, we fine-tune both the LLM backbone and projector parameters, with all experiments performed using full parameter fine-tuning. For InternVL2.5, we use the fine-tuning scripts provided in their official documentation [^1] https://internvl.readthedocs.io/en/latest/internvl2.5/finetune.html. For LLaVA-OV-7B, we perform fine-tuning using the Swift (Zhao et al., 2024) framework. All experiments are conducted on 8 A100-80G GPUs.
@@
For Textual SFT, we follow the settings of Hu et al. (2024), sampling 2k harmful and 1k benign samples from SafeRLHF Ji et al. (2024). For the harmful input data, we use Llama3-8B-Instruct Inan et al. (2023) to generate safe responses, while the safe samples use the original responses from the dataset. Since the data is in the text domain, following Hu et al. (2024), we match each sample with an all-black image during fine-tuning. It is worth noting that for Qwen2-VL, we directly use the open-sourced fine-tuned model [^2] https://huggingface.co/Foreshhh/Qwen2-VL-7B-SafeRLHF from Hu et al. (2024), while for the recent InternVL2.5, the detailed training parameters are reported in Table 7.
@@
VLGuard-M mixes VLGuard training data into general data during the instruction tuning stage. Given that we cannot access the training data for Qwen2-VL and InternVL2.5, we can only conduct experiments on VLGuard Posthoc Fine-Tuning Zong et al. (2024), which we refer to as VLGuard-P. This includes 2k multimodal single-image safety-related data and 1k benign samples. As for VLGuard-R, mentioned in Sec. 2, we designed a prompt that guides InternVL2.5-78B to first analyze the input content and provide reasons for its potential unsafe nature, ultimately generating harmless responses. This process was used to construct SFT labels for 2k unsafe inputs. For the 1k safe samples, we directly use the original responses from VLGuard. The specific prompt is reported in Figure. As with Textual SFT, for Qwen2-VL, we use the open-sourced VLGuard-P [^3] https://huggingface.co/Foreshhh/Qwen2-VL-7B-VLGuard from Hu et al. (2024). The training details for VLGuard-R and InternVL2.5 are provided in Table 7. Additionally, the detailed SFT label construction prompt is shown in Appendix F.2.
@@
With our proposed MIS training set, which contains 4k safety-related multi-image samples, we introduce MIRage. Given that safety risks in MIS arise from text instructions linking objects across two images, we design a prompt to guide InternVL2.5-78B to respond based on safety reasoning logic. Specifically, as shown in the figure, the model is first instructed to identify the elements in both images, then reason about the potential risks created by these elements under the text instruction, and finally provide a safe response. Building on this, we additionally include 500 general QA samples to ensure that the model’s instruction-following capability remains unaffected. The detailed training parameters are provided in Table 7. Additionally, the detailed SFT label construction prompt is shown in Appendix F.2.
B.3 Evaluation Details
B.3.1 Evaluation on MIS